OpenStudy (anonymous):

A metal washer is made by cutting out a circle 13.2 mm in diameter from a larger circle. To the nearest tenth, what is the area of the washer? larger is 4.4mm

OpenStudy (anonymous):

A metal washer is made by cutting out a circle 13.2 mm in diameter from a larger circle. Larger Circle is 4.4mm. To the nearest tenth, what is the area of the washer? 243.3 mm2 60.8 mm2 547.1 mm2 972.6 mm2 I got 60.8mm2 is this correct

OpenStudy (mrnood):

How can you cut a circle of 13.2dia from a disc which is only 4.4? Please check your question

OpenStudy (mrnood):

 Area of Large circle = pi R^2 Area of small circle = pi r^2 Area of washer = pi (R^2-r^2)